Assume a closed economy and no government. Also assume consumption
C=50-0.8Yd and investment I=80
derive the equation for saving
1
Expert's answer
2020-06-18T13:47:37-0400

"S=-a+(1-b)(Y-T)"

Given ,C=50-0.8Yd. Therefore,a=50 and b= 0.8.Since there is no government intervention T=0.

S=-50+(1-0.8)(Y-0)

S=-50+0.2Y